# The StudyChat Dataset: Exploring Student Dialogues With ChatGPT in an Artificial Intelligence Course

### v3

@@ 4.3 Clustering Students by Behavior
To investigate whether distinct patterns of student behavior emerge, we apply unsupervised clustering to behavioral interaction features. We follow an approach similar to Brender et al. (2024), forming feature vectors that normalize each student’s DA utterance types by their total count of utterances. In total, there are 39 features, with the 8 broad DA features and 31 specific features. These vectors represent each student’s distribution of DA labels over the course of the semester. We use K-means clustering on these feature vectors to group students based on their behavioral profiles, applying the method separately to Fall and Spring semester data. The optimal number of clusters, $k=4$, is determined using the elbow method and silhouette analysis.
@@
We report the distribution of DA for each cluster across for the fall and spring in Table 5 and  6 respectively. While the exact makeup of each cluster varies across semesters, some general patterns emerge. For example, certain broad labels, such as writing request and contextual question labels, are prevalent across multiple clusters, while others (e.g., off-topic or conceptual questions) are concentrated in specific clusters. Notably, some clusters represent high engagement (e.g., students with over 100 total utterances), while others capture minimal or off-topic usage. These differences suggest a spectrum of interaction styles, from consistent academic engagement to sparse or misaligned LLM usage. We see two recurring patterns of cluster assignment: groups with frequent writing requests and groups with frequent question-asking requests.
@@
We examine the differences in the similar clusters in the fall semester and find patterns emerge. For the writing-request clusters, students in cluster 0 tend to ask the chatbot to write code while students in cluster 3 display a mixture of code writing and english writing requests. This observation suggests that students in cluster 3 lean on the LLM not only for help in implementing the requested program, but also in synthesizing the results and writing assignment reports. For the question-asking clusters, students in cluster 1 tend to ask conceptual questions, mostly on and Python-related questions and computer science topics, while students in cluster 2 ask a mixture of conceptual questions and contextual questions of a much wider spread of topics. This observation suggests that students in cluster 1 use the LLM for understanding the tools needed to complete the assignments, while students in the later category use the LLM as a more general purpose tutoring system. In examining the average assignment and exam scores between these clusters, we see that all clusters have comparable average assignment scores. However coding-question groups trends towards a higher average score on the exams $89\pm 7.3$ while report-writers tend to have a marginally lower average score $83.8\pm 11.7$. These observations suggest that students who ask targeted questions of the LLM for specific programming and coding concepts while working on assignments tend to enter exams with a better understanding of the material than students who request extensive writing assistance, which may circumvent the learning objectives of the assignments.
@@
To further explore these trends in assessment scores, we again perform a regression analysis. We use the same baseline model, a single independent variable, running average of prior assessment scores, to dependent variable of assessment grades. Similar to the usage interaction analysis, for cluster membership we include four additional binary variables, each associated with a group to assess its additional explanatory power. We find that adding the cluster variable improved the predictive performance marginally. For the Fall semester, the mean $R^{2}$ increased from 0.1783 to 0.2152; for Spring, it improved from 0.1682 to 0.1844. We conduct F-tests for each regression and find that no indications of statistically significant improvement in explanatory power, with the highest Resulting F-score of F=2.67, p=0.06. These results demonstrate that while this behavioral clustering does provide an improvement, similar to the usage grouping, we cannot definitively conclude this clustering strategy is statistically meaningful in terms of understanding course outcomes. However, while the clusters are not definitive predictors, they do offer interpretable groupings of student behavior that correlate with distinct behavior styles ranging from frequent, academically aligned interactions to sparse or off-topic usage. This diversity highlights that engagement is multidimensional, varying not only in frequency, but also type of behavior. These clusters provide quantitative insights into patterns of student LLM usage which may prove helpful in educating students to use these tools in a learning-positive way.
5 Conclusions, Limitations, Future Work
@@
In this work, we present the StudyChat dataset, a collection of over 16,000 interactions between students and ChatGPT in a university-level artificial intelligence course. We systematically create and apply a dialogue act labeling schema applied to this dataset via LLM labeling, validate this labeling with human agreement, and conduct a regression and clustering analysis to analyze trends of student behavior as they relate to course outcomes.
@@
While our analysis offers new insights into how students interact with LLMs, the takeaways from our work have limitations. First, our dataset was drawn from a single upper-division AI course, which limits how well the findings generalize. Second, participants in this study were aware that their conversations were recorded and thus their behaviors may have been altered and subject to the Hawthorne Effect landsberger1958hawthorne.
@@
Our findings from this exploration and the resulting dataset open many avenues for future work. First, since the StudyChats dataset also includes student assignment submissions, we can further compare these submissions against student utterances when conversing with the LLM which may reveal further behavioral insights and trends. Second, our initial labeling schema and labeling methodology could be refined further, to better cover overlapping labels and missing dialogue acts with more accurate and fine-grained methods of classifying student utterances. Finally, these insights can be taken back into a classroom setting where we could create LLM-based tutoring tools that detect and discourage behaviors associated with negative course outcomes.
Acknowledgments
@@
We would like to thank all the students of this class for their participation and engagement in this project. We are also grateful to the entire course staff team for their hard work and dedication to teaching.
We extend a special thanks to Aditya Singh for his insightful discussions which helped shape this work, and for his contributions to the design and implementation of many course assignments.
@@ Appendix A Appendix
Table 3: Distribution of dialogue act labels in StudyChat.
@@
Table 4: Summary of statistically significant regression coefficients ($p<0.05$) for predicting assessment outcomes from interaction features, grouped by semester and feature set. Only significant results are listed.
Semester Feature Set Feature Name Target Coefficient $p$ $n$
F24 Broad No significant coefficients
Specific Contextual Questions: Code Explanation Assignment 3 0.0065-0.0065 $0.021$ $72$
Contextual Questions: Other Assignment 3 0.0156-0.0156 $0.034$ $72$
Off Topic: Chit-Chat Assignment 3 $0.0235$ $0.043$ $72$
Provide Context: Code Assignment 4 0.0426-0.0426 $0.001$ $66$
Verification: Other Assignment 4 0.0291-0.0291 $0.007$ $66$
Conceptual Questions: Other Concept Assignment 5 0.0575-0.0575 $0.025$ $64$
Writing Request: Other Assignment 5 0.0590-0.0590 $0.033$ $64$
S25 Broad Conceptual Questions Exam 1 $0.0133$ $0.024$ $104$
Editing Request Exam 1 $0.0296$ $0.017$ $104$
Contextual Questions Assignment 5 0.0020-0.0020 $0.005$ $72$
Specific Conceptual Questions: Python Library Exam 1 $0.0234$ $0.010$ $104$
Contextual Questions: Assignment Clarification Assignment 3 0.0034-0.0034 $0.021$ $103$
Conceptual Questions: Other Concept Assignment 3 $0.0095$ $0.029$ $103$
Writing Request: Conversion Exam 2 $0.0028$ $0.045$ $119$
Contextual Questions: Assignment Clarification Assignment 5 0.0072-0.0072 $0.039$ $72$
Contextual Questions: Code Explanation Assignment 5 $0.0009$ $0.044$ $72$
Contextual Questions: Other Assignment 5 0.0075-0.0075 $0.017$ $72$
Conceptual Questions: Other Concept Assignment 5 $0.0016$ $0.019$ $72$
Conceptual Questions: Mathematics Assignment 6 0.1062-0.1062 $0.002$ $74$
Off Topic: Gratitude Exam 3 0.0337-0.0337 $0.025$ $119$
Figure 4: Usage over time organized by assignments for the Fall 2024 semester
Figure 5: Usage over time organized by assignments for the Spring 2025 semester
Figure 6: Box plot showing the distribution of total utterances per student for each conversation topic where outliers are shown with cross markers.
Table 5: Top 10 Dialogue Acts per Cluster – F24 Semester
@@
Table 6: Top 10 Dialogue Acts per Cluster – S25 Semester

### v4

@@ 4.3. Clustering Students by Behavior
To investigate whether distinct patterns of student behavior emerge, we apply unsupervised clustering to behavioral interaction features. We follow an approach similar to (Brender et al., 2024), forming feature vectors that normalize each student’s DA utterance types by their total count of utterances. In total, there are 39 features, with the 8 broad DA features and 31 specific features. These vectors represent each student’s distribution of DA labels over the entire semester. We use K-means clustering on these feature vectors to group students based on their behavioral profiles, applying the method separately to Fall and Spring semester data. The optimal number of clusters, $k=4$, is determined using the elbow method and silhouette analysis.
@@
We report the distribution of the most common DA for each cluster in the table at the bottom of Figure 4. While the exact makeup of each cluster varies across semesters, some general patterns emerge. For example, certain broad labels, such as writing request and contextual question labels, are prevalent across multiple clusters, while others (e.g., off-topic or conceptual questions) are concentrated in specific clusters. Notably, some clusters represent high engagement (e.g., students with over 100 total utterances), while others capture minimal or off-topic usage. These differences suggest a spectrum of interaction styles, from consistent academic engagement to sparse or misaligned LLM usage. We see two recurring patterns of cluster assignment: groups with frequent writing requests and groups with frequent question-asking requests.
# Cluster n Exam Assign.
0 Code Writers 25 $87.0\pm 0.6$ $95.6\pm 2.5$
1 Coding Question 16 $89.0\pm 7.3$ $96.1\pm 2.1$
2 General Question 27 $86.5\pm 8.0$ $94.4\pm 6.3$
3 Report Writers 16 $83.8\pm 11.7$ $93.7\pm 4.4$
Figure 5. Comparison of dialogue act PCA clusters from Fall 2024. The radar plot (top) shows the top 10 most prominent specific dialog acts with values indicating the mean percentage of that act type within each cluster. The table (bottom) reports the number of students in each cluster and their mean aggregate course scores. Exam scores of cluster 3 are lower on average than other clusters.
@@
We visualize the most predominant DAs and student grades for the fall semester PCA clusters in Figure 5. We observe further patterns in examining the behavior patterns and course outcomes in these clusters. For the writing-request clusters, students in cluster 0 tend to ask the chatbot to write code, while students in cluster 3 display a mixture of code writing and English writing requests. This observation suggests that students in cluster 3 lean on the LLM not only for help in implementing the requested program but also in synthesizing the results and writing assignment reports. For the question-asking clusters, students in cluster 1 tend to ask conceptual questions, mostly on Python-related questions and computer science topics, while students in cluster 2 ask a mixture of conceptual questions and contextual questions of a much wider spread of topics. This observation suggests that students in cluster 1 use the LLM for understanding the tools needed to complete the assignments, while students in the latter category use the LLM as a more general-purpose tutoring system. In examining the average assignment and exam scores between these clusters, we see that all clusters have comparable average assignment scores. However, coding-question groups trend towards a higher average score on the exams $89\pm 7.3$ while report-writers tend to have a marginally lower average score $83.8\pm 11.7$. These observations suggest that students who ask targeted questions of the LLM for specific programming and coding concepts while working on assignments tend to enter exams with a better understanding of the material than students who request extensive writing assistance, which may circumvent the learning objectives of the assignments.
@@
To further explore these trends in assessment scores, we again perform a regression analysis. We use the same baseline model, a single independent variable, running average of prior assessment scores, to dependent variable of assessment grades. Similar to the usage interaction analysis, for cluster membership, we include four additional binary variables, each associated with a group to assess its additional explanatory power. We find that adding the cluster variable improved the predictive performance marginally. For the Fall semester, the mean $R^{2}$ increased from 0.1783 to 0.2152; for Spring, it improved from 0.1682 to 0.1844. We conduct F-tests for each regression and find no indications of statistically significant improvement in explanatory power, with the highest resulting F-score of F=2.67, p=0.06. These results demonstrate that while this behavioral clustering does provide an improvement, similar to the usage grouping, we cannot definitively conclude that this clustering strategy is statistically meaningful in terms of understanding course outcomes. However, while the clusters are not definitive predictors, they do offer interpretable groupings of student behavior that correlate with distinct behavior styles ranging from frequent, academically aligned interactions to sparse or off-topic usage. While our clustering analysis reveals some behavior types that are associated with lower exam performance, further work is needed to achieve consistent and reliable detection of behaviors that are misaligned with learning objectives.
5. Conclusions, Limitations, Future Work
@@
In this work, we present the StudyChat dataset, a collection of over 16,000 interactions between students and ChatGPT in a university-level artificial intelligence course. We systematically create and apply a dialogue act labeling schema applied to this dataset via LLM labeling, validate this labeling with human agreement, and conduct regression and clustering analysis to analyze trends of student behavior as they relate to course outcomes.
@@
While our analysis offers new insights into how students interact with LLMs, the takeaways from our work have limitations. First, our dataset was drawn from a single upper-division AI course, which limits how well the findings generalize. Second, participants in this study were aware that their conversations were recorded and thus their behaviors may have been altered and subject to the Hawthorne Effect (Landsberger, 1958).
@@
Our findings from this exploration and the resulting dataset open many avenues for future work. First, since the StudyChats dataset also includes student assignment submissions, we can further compare these submissions against student utterances when conversing with the LLM, which may reveal further behavioral insights and trends. Second, our initial labeling schema and labeling methodology could be refined further to better cover overlapping labels and missing dialogue acts with more accurate and fine-grained methods of classifying student utterances. Finally, these insights can be taken back into a classroom setting where we could create LLM-based tutoring tools that detect and discourage behaviors associated with negative course outcomes.
Acknowledgements.
@@
We would like to thank all the students of this course for their participation and engagement in this project. We are also grateful to the entire course staff for their hard work and dedication to teaching. We extend a special thanks to Aditya Singh for his contributions to the design and implementation of many course assignments, as well as the insightful discussions he shared with the authors regarding this work. The authors are partially supported by the NSF under grants 2237676 and 2418657.
@@ Appendix A Appendix
Table 3. Distribution of dialogue act labels in StudyChat.
@@
Table 4. Top 10 Dialogue Acts per Cluster – F24 Semester
@@
Table 5. Top 10 Dialogue Acts per Cluster – S25 Semester
@@
Figure 6. Usage over time organized by assignments for the Fall 2024 semester
Figure 7. Usage over time organized by assignments for the Spring 2025 semester
Figure 8. Box plot showing the distribution of total utterances per student for each conversation topic where outliers are shown with cross markers.